Solve equations by factoring
Solve the equation.
(x + 8)(x - 4) = -20

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
(x + 8)(x - 4) = -20
x^2 +4x -32 = -20
x^2 + 4x -12 = 0
(x+6)(x-2) = 0 Note:SUM of the inner product(6x) and the outer product(-2x) = 4x
(x+6)=0
x = -6
(x-2) = 0
x = 2